OpenStudy (anonymous):

Find all roots of the equation z^2-j=0

OpenStudy (kainui):

Any guesses? What's the root of an equation?

OpenStudy (anonymous):

z-1

OpenStudy (anonymous):

I understand that it must be in the form of ax^2+bx+c=0 So does a=1, b=-1 and c=0?

OpenStudy (kainui):

Nope, it's: a=1, b=0, c=-j

OpenStudy (anonymous):

so then that is plugged into the quadratic equation?

OpenStudy (anonymous):

sorry I meant to write: \[\frac{ 2\sqrt{1j} }{ 2 }\]